The legs of a right triangle have lengths of 12 inches and 17 inches what is the length of its hypotenuse to the nearest tenth of an inch?
Answer by ewatrrr(24785)

Applying the Pythagorean Theorem
hyp = sqrt(144 + 289) = sqrt(433) = 20.8in (nearest tenth)